# Fooling the Textual Fooler via Randomizing Latent Representations

### v1

@@ Abstract
Despite outstanding performance in a variety of NLP tasks, recent studies have revealed that NLP models are vulnerable to adversarial attacks that slightly perturb the input to cause the models to misbehave. Among these attacks, adversarial word-level perturbations are well-studied and effective attack strategies. Since these attacks work in black-box settings, they do not require access to the model architecture or model parameters and thus can be detrimental to existing NLP applications. To perform an attack, the adversary queries the victim model many times to determine the most important words in an input text and to replace these words with their corresponding synonyms. In this work, we propose a lightweight and attack-agnostic defense whose main goal is to perplex the process of generating an adversarial example in these query-based black-box attacks; that is to fool the textual fooler. This defense, named AdvFooler, works by randomizing the latent representation of the input at inference time. Different from existing defenses, AdvFooler does not necessitate additional computational overhead during training nor relies on assumptions about the potential adversarial perturbation set while having a negligible impact on the model’s accuracy. Our theoretical and empirical analyses highlight the significance of robustness resulting from confusing the adversary via randomizing the latent space, as well as the impact of randomization on clean accuracy. Finally, we empirically demonstrate near state-of-the-art robustness of AdvFooler against representative adversarial word-level attacks on two benchmark datasets.
@@ 1 Introduction
To generate adversarial examples, adversarial attack methods manipulate different aspects of the input sentence, from introducing character errors such as typos or visually similar characters (Gao et al., 2018a; Eger et al., 2019) and replacing words without significantly changing the original semantic in the perturbed inputs (Jin et al., 2020; Li et al., 2020) to recreating sentences with similar meanings using paraphrasing (Iyyer et al., 2018; Qi et al., 2021). Among these methods, adversarial-word substitutions (Alzantot et al., 2018) are some of the most widely studied and effective adversarial approaches. These methods query the victim model multiple times to find the corresponding adversarial perturbations of important words of an input text.
@@
In response to the adversarial threat and to enhance the robustness of NLP models, numerous defenses against textual adversarial attacks have been developed (Zhu et al., 2020; Si et al., 2021; Wang et al., 2020; Ye et al., 2020; Shi et al., 2020b; Xu et al., 2020). The most popular approaches are adversarial training (Zhu et al., 2020; Dong et al., 2021; Madry et al., 2017), which augments the training data with adversarial examples using an additional optimization step, and randomized smoothing (Ye et al., 2020; Zeng et al., 2021), which replaces the model with its stochastic ensemble based on random perturbations of the input; both of these approaches require training modifications and additional non-trivial training computation (e.g., for adversarial training). Due to the discrete nature of texts, to create adversarial inputs, these defenses also substitute input words with adversarial words sampled from predefined perturbation sets constructed based on the potential attacks instead of gradient-based optimization. This substitution process also aims to preserve the original input semantics. Nonetheless, assuming knowledge of the potential attacks is often unrealistic and impractical.
@@
In this paper, we propose a lightweight, attack-agnostic defensive method that can increase the robustness of NLP models against textual adversarial attacks. Our defense, called AdvFooler, randomizes the latent representation of the input at test time to fool the adversary throughout their attack, which typically involves iteratively sampling of discrete perturbations to generate an adversarial sample. Being a test-time defense with negligible computational overhead, AdvFooler also does not incur any training-time computation. Furthermore, as AdvFooler operates within the latent space of the model, it remains agnostic of the perturbation sets associated with potential attacks. While there are some randomization defenses in NLP (Ye et al., 2020; Zeng et al., 2021), they rely on randomized smoothing with necessary modifications to the model’s training to reduce the variance in the model’s outputs caused by randomizing the input or embedding space. The advantages of AdvFooler over the other defenses are shown in Table 1. Our contributions can be summarized as follows:
@@
  • We extensively evaluate AdvFooler through experiments on various benchmark datasets and representative attacks. The results demonstrate that AdvFooler is on par with state-of-the-art defenses while requiring few modeling assumptions, being pluggable, and incurring negligible additional computational overhead.
@@ 2.1 Adversarial attacks in NLP
Adversarial attacks in NLP can be divided into two categories, depending on what information the attacker has on the victim model. These categories are white-box attacks (Ebrahimi et al., 2018; Cheng et al., 2020) - those that require access to the model’s architecture and its parameters - and black-box attacks (Li et al., 2020; Jin et al., 2020; Li et al., 2018) - those that rely only querying the model for its output. Since the model user rarely shares the underlying model’s architecture of their NLP service, let alone their model’s parameters, white-box attacks have extremely limited practicality, compared to black-box attacks. In the paper, we will focus on the black-box, query-based attack setting, as it is a more realistic scenario in practice and is commonly studied in similar works (Zeng et al., 2021; Nguyen Minh & Luu, 2022; Zhang et al., 2022). Note that, there are also black-box, transfer-based attacks, which, however, still require knowledge of the model architecture for the attack to be effective (Yuan et al., 2021; Li et al., 2021a).
@@
where $d(x_{adv}-x)$ is the perceptual distance between $x$ and $x_{adv}$, and $\sigma$ is the maximum acceptable distance. The distance function $d$ can measure the number of perturbed words between the original and adversarial input (Gao et al., 2018b). Another popular distance is the semantic dissimilarity between a pair of texts (Iyyer et al., 2018; Li et al., 2020) measured as the cosine distance between the corresponding embedding vectors of the inputs extracted from Universal Sentence Encoder (USE) (Cer et al., 2018).
2.2 Adversarial word substitution
Among textual adversarial attacks, adversarial word substitution is one of the most widely studied approaches. These attacks create an adversarial example of an input text $x$ by first identifying the most important words in $x$ and then replacing those words with synonyms from their corresponding synonym sets to cause the victim model to misbehave while aiming to preserve the original semantic meaning. The synonym sets can be created using word embedding models (Mrkšić et al., 2016; Pennington et al., 2014) or using the predictions of a PrLMs given the input sentence (Li et al., 2020). One important feature of these attacks is that both of the aforementioned steps require querying the victim model many times to determine which words are important or which synonyms maximize the model’s prediction errors. In the first step, the adversary queries the model to calculate the importance score of each word in $x$ and select the most potential locations (with the highest importance scores) for later perturbations. To find the substituting synonyms that maximize the model’s prediction error, the attacks can either perform a greedy search (Li et al., 2020; Jin et al., 2020; Li et al., 2018) or combinatorial optimization algorithm (Zang et al., 2020; Alzantot et al., 2018). Similarly, this step involves a sequence of queries on the victim model. By introducing randomization to the latent representations of queries, our defensive mechanism disrupts the adversary’s ability to estimate important words, making their synonym-substituting process significantly harder during the attack.
@@ 2.3 Adversarial defense methods
To defend against textual adversarial attacks, several defenses (Wang et al., 2022) have been developed. The goal of an adversarial defense is to improve the robustness of the victim model; that is to achieve good performance on both clean and adversarial examples. Existing adversarial defenses can be divided into two categories: certified (Ye et al., 2020; Zeng et al., 2021; Jia et al., 2019) and empirical (Zhu et al., 2020; Wang et al., 2020; Dong et al., 2021; Li & Qiu, 2020; Zhou et al., 2021; Miyato et al., 2016). First introduced in Goodfellow et al. (2014), adversarial training is the most often studied empirical defense method. It provides additional regularization to the model and improves the model’s robustness when training it with adversarial samples. Several subsequent works study adversarial training for NLP tasks with adversarial examples created on the input space (Ren et al., 2019; Li et al., 2020; Jin et al., 2020). Some adversarial training works improve the model’s robustness by introducing adversarial perturbations in the embedding space (Zhu et al., 2020), or by incorporating inductive bias to prevent the model from learning spurious correlations (Wang et al., 2020; Madry et al., 2017). These methods effectively improve the model’s robustness without significant compromise to its clean accuracy.
In contrast to the empirical methods, certified defenses can provably guarantee model robustness even under sophisticated attackers. One popular certified-defense approach is randomized smoothing (Ye et al., 2020; Zeng et al., 2021), which constructs a set of stochastic ensembles from the input and leverages their statistical properties to provably guarantee robustness. In contrast, Interval Bound Propagation (IBP) methods (Shi et al., 2020a; Jia et al., 2019; Huang et al., 2019) utilizes axis-aligned bounds to confine adversarial examples. Despite their effectiveness and ability to certify robustness, both approaches rely on access to the synonym sets of potential attacks. Randomized-smoothing methods also modify the training phase to reduce the variance of the outputs of the ensemble. Due to the ensemble, they incur significant computational overhead in the inference phase. In contrast, our defense is lightweight and attack-agnostic, and can better control the variance of the outputs by randomizing the latent representations, instead of via randomization in the input space. A summary of the characteristics of AdvFooler and the existing defenses is shown in Table 1.
@@ 3.1 Randomized latent-space defense against adversarial word substitution
Since the adversary relies on querying the victim model multiple times to find adversarial examples, receiving fluctuating feedback from the model will make it significantly harder for the adversary to find the optimal adversarial perturbations. To fool such an adversary, we introduce stochasticity to the model by randomizing the latent representation of an input $x$. Formally, for $h_{l}$ as the $l-$th layer of the model, we sample an independent noise vector $\epsilon$, which is added to $h_{l}(x)$ as input to the next layer of the model. Without loss of generality, $\epsilon$ is sampled from a Gaussian distribution $\mathcal{N}(0,\Sigma)$ with $\Sigma$ as a diagonal covariance matrix, or $\mathcal{N}(0,\nu I),\nu\in\mathcal{R}$. The detailed algorithm is presented in Algorithm 1.
Let $f_{\operatorname{AdvFooler}}$ be the proposed randomized model corresponding to the original $f$. When the variance of injected noise is low, we can assume that small noise slightly changes the output logits but does not shift the prediction of the model. In other words, the mean of the randomized model $f_{\operatorname{AdvFooler}}$ with input $x$ is exactly the prediction of $f$ for $x$. Consequently, adversarial samples of $f$ are adversarial samples of $f_{\operatorname{AdvFooler}}$.
Algorithm 1 AdvFooler’s Randomized Latent Defense
Input: a model $f$, an input sentence $x$, noise magnitude $\nu$.
@@
6:return $z_{L}$
@@ 3.2 Effect of randomizing latent representations on adversarial attacks
In textual adversarial attacks, the adversary first determines important words in a sentence $x=\{w_{0},\dots,w_{i}\}$ using the importance score, which is computed for each word $w_{i}$ as
$I_{w_{i}}=f_{y}(x)-f_{y}(x_{/w_{i}})$ (1)
where $f_{y}$ is the logit returned by the model $f$ w.r.t ground-truth label $y$. The adversary then selects words with the highest scores to perturb. Intuitively, this process emulates estimating the gradient of the importance score in the discrete space.
When introducing randomness to the latent space, the important scores are changed. Such changes mislead the attack to select a different set of words to perturb. Perturbing these “less important” words (those that have lower original importance scores before randomization) makes the synonym replacement phase significantly more challenging for the attack, thus reducing the attack’s success probability.
Theorem 1.
[…]
Theorem 1 (proof is in the supplementary material) states that randomly perturbing the hidden presentation leads to a randomized important score. When $\nu$ is high, the randomized importance scores have high variance, which can cause the attack to select the wrong important words.
@@ 3.3 Effect of randomizing latent representations on clean accuracy
If randomization induces substantial fluctuation in the model’s output, there is a possibility that the model might predict a wrong label, resulting in a decrease in clean accuracy. Randomized defenses such as SAFER (Ye et al., 2020) and RanMASK (Zeng et al., 2021) introduce randomness to the model in the input space (i.e., randomizing the input tokens). While they are also effective against textual adversarial attacks, the input space of NLP models is discrete. Perturbing discrete inputs can lead to significant changes in the model’s outputs (a fact that we will empirically prove in the next section), which results in a different model’s prediction. To mitigate this problem, existing randomized input defenses rely on randomized smoothing: train the model with perturbed input and apply ensemble at inference to reduce the variance in the model’s output. Nevertheless, these approaches are computationally expensive and require access to the training phase. Conversely, by randomizing the latent representations, AdvFooler induces a significantly smaller variance in the model’s output. Furthermore, AdvFooler gives the defender flexibility to select the suitable induced variance without requiring access to the training process. Specifically, given a pretrained classification model and a small clean test set, the defender can select the noise scale $\nu$ at which the variance causes the clean accuracy drops by a chosen percentage (e.g., in our experiments, it is 1%).

### v2

@@ Abstract
Despite outstanding performance in a variety of Natural Language Processing (NLP) tasks, recent studies have revealed that NLP models are vulnerable to adversarial attacks that slightly perturb the input to cause the models to misbehave. Several attacks can even compromise the model without requiring access to the model architecture or model parameters (i.e., a black-box setting), and thus are detrimental to existing NLP applications. To perform these attacks, the adversary queries the victim model many times to determine the most important parts in an input text and transform. In this work, we propose a lightweight and attack-agnostic defense whose main goal is to perplex the process of generating an adversarial example in these query-based black-box attacks; that is to fool the textual fooler. This defense, named AdvFooler, works by randomizing the latent representation of the input at inference time. Different from existing defenses, AdvFooler does not necessitate additional computational overhead during training nor does it rely on assumptions about the potential adversarial perturbation set while having a negligible impact on the model’s accuracy. Our theoretical and empirical analyses highlight the significance of robustness resulting from confusing the adversary via randomizing the latent space, as well as the impact of randomization on clean accuracy. Finally, we empirically demonstrate near state-of-the-art robustness of AdvFooler against representative adversarial attacks on two benchmark datasets. Code for AdvFooler is available at: https://github.com/mail-research/AdvFooler-text-defender
**footnotetext: The work was done prior to joining Amazon.
@@ 1 Introduction
To generate adversarial examples, adversarial attack methods manipulate different aspects of the input sentence, from introducing character errors such as typos or visually similar characters (Gao et al., 2018a; Eger et al., 2019) and replacing words without significantly changing the original semantic in the perturbed inputs (Jin et al., 2020; Li et al., 2020) to recreating sentences with similar meanings using paraphrasing (Iyyer et al., 2018; Qi et al., 2021). Among these methods, black-box adversarial attacks (Alzantot et al., 2018) are some of the most widely studied and effective adversarial approaches. These methods query the victim model multiple times to find the corresponding adversarial perturbations of important words of an input text.
@@
In response to the adversarial threat and to enhance the robustness of NLP models, numerous defenses against textual adversarial attacks have been developed (Zhu et al., 2020; Si et al., 2021; Wang et al., 2021; Ye et al., 2020; Shi et al., 2020a; Xu et al., 2020). The most popular approaches are adversarial training (Zhu et al., 2020; Dong et al., 2021; Madry et al., 2018), which augments the training data with adversarial examples using an additional optimization step, and randomized smoothing (Ye et al., 2020; Zeng et al., 2023), which replaces the model with its stochastic ensemble based on random perturbations of the input; both of these approaches require training modifications and additional non-trivial training computation (e.g., for adversarial training). Due to the discrete nature of texts, to create adversarial inputs, these defenses also substitute input words with adversarial words sampled from predefined perturbation sets constructed based on the potential attacks instead of gradient-based optimization. This substitution process also aims to preserve the original input semantics. Nonetheless, assuming knowledge of the potential attacks is often unrealistic and impractical.
@@
In this paper, we propose a lightweight, attack-agnostic defensive method that can increase the robustness of NLP models against textual adversarial attacks. Our defense, called AdvFooler, randomizes the latent representation of the input at test time to fool the adversary throughout their attack, which typically involves iteratively sampling of discrete perturbations to generate an adversarial sample. Being a test-time defense with negligible computational overhead, AdvFooler also does not incur any training-time computation. Furthermore, as AdvFooler operates within the latent space of the model, it remains agnostic of the perturbation sets associated with potential attacks. While there are some randomization defenses in NLP (Ye et al., 2020; Zeng et al., 2023), they rely on randomized smoothing with necessary modifications to the model’s training to reduce the variance in the model’s outputs caused by randomizing the input or embedding space. The advantages of AdvFooler over the other defenses are shown in Table 1. Our contributions can be summarized as follows:
@@
  • We extensively evaluate AdvFooler through experiments on various benchmark datasets and representative attacks. The results demonstrate that AdvFooler is a competitive defense, compared to the existing representative textual adversarial defenses, while being under more constraints, including few modeling assumptions, being pluggable, and incurring negligible additional computational overhead.
@@ 2.1 Adversarial attacks in NLP
Adversarial attacks in NLP can be divided into two categories, depending on what information the attacker has on the victim model. These categories are white-box attacks (Ebrahimi et al., 2018; Cheng et al., 2020) - those that require access to the model’s architecture and its parameters - and black-box attacks (Li et al., 2020; Jin et al., 2020; Li et al., 2018) - those that rely only querying the model for its output. Since the model user rarely shares the underlying model’s architecture of their NLP service, let alone their model’s parameters, white-box attacks have extremely limited practicality, compared to black-box attacks. In the paper, we will focus on the black-box, query-based attack setting, as it is a more realistic scenario in practice and is commonly studied in similar works (Zeng et al., 2023; Nguyen Minh and Luu, 2022; Zhang et al., 2022). Note that, there are also black-box, transfer-based attacks, which, however, still require knowledge of the model architecture for the attack to be effective (Yuan et al., 2021; Li et al., 2021a).
@@
where $d(x_{adv}-x)$ is the perceptual distance between $x$ and $x_{adv}$, and $\sigma$ is the maximum acceptable distance. For example, $d$ can measure the number of perturbed words between the original and adversarial input (Gao et al., 2018b). Another popular distance is the semantic dissimilarity between a pair of texts (Iyyer et al., 2018; Li et al., 2020), defined as the cosine distance between the extracted embedding vectors of the inputs from Universal Sentence Encoder (USE) (Cer et al., 2018).
Solving the objective in Eq. (2.1) is equivalent to searching for perceptually valid perturbations that flip the prediction of the victim model. However, this is a challenging task due to the considerably large search space for textual data. Specifically, the works in Mrkšić et al. (2016); Li et al. (2018) demonstrate that perturbing random words in the sentence fails to effectively fool the model with a reasonable number of queries. To accelerate the search procedure, existing attacks locate important parts of the sentence, e.g., word Li et al. (2018); Jin et al. (2020); Li et al. (2020) or group of words Chen et al. (2021); Lei et al. (2022), to perturb.
Among textual adversarial attacks, adversarial word substitution is one of the most widely studied approaches. These attacks create an adversarial example of an input text $x$ by first identifying the most important words in $x$ and then replacing those words with synonyms from their corresponding synonym sets to cause the victim model to misbehave while aiming to preserve the original semantic meaning. The synonym sets can be created using word embedding models (Mrkšić et al., 2016; Pennington et al., 2014) or using the predictions of a PrLMs given the input sentence (Li et al., 2020). One important feature of these attacks is that both of the aforementioned steps require querying the victim model many times to determine which words are important or which synonyms maximize the model’s prediction errors. In the first step, the adversary queries the model to calculate the importance score of each word in $x$ and select the most potential locations (with the highest importance scores) for later perturbations. To find the substituting synonyms that maximize the model’s prediction error, the attacks can either perform a greedy search (Li et al., 2020; Jin et al., 2020; Li et al., 2018) or combinatorial optimization algorithm (Zang et al., 2020; Alzantot et al., 2018). Chen et al. (2021) extend this principle to multi-granularity by determining important constituents in the syntactic tree that maximizes the decrease in confidence score when being paraphrased. Similarly, this step involves a sequence of queries on the victim model. By introducing randomization to the latent representations of queries, our defensive mechanism disrupts the adversary’s ability to estimate important words (or paraphrases), making their synonym-substituting process significantly harder during the attack.
@@ 2.2 Adversarial defense methods
To defend against textual adversarial attacks, several defenses (Wang et al., 2022) have been developed. The goal of an adversarial defense is to improve the robustness of the victim model; that is to achieve good performance on both clean and adversarial examples. Existing adversarial defenses can be divided into two categories: certified (Ye et al., 2020; Zeng et al., 2023; Jia et al., 2019) and empirical (Zhu et al., 2020; Wang et al., 2021; Dong et al., 2021; Li and Qiu, 2020; Zhou et al., 2021; Miyato et al., 2017).
First introduced in Goodfellow et al. (2014), adversarial training is the most often studied empirical defense method. It provides additional regularization to the model and improves the model’s robustness when training it with adversarial samples. Several subsequent works study adversarial training for NLP tasks with adversarial examples created on the input space (Ren et al., 2019; Li et al., 2020; Jin et al., 2020). Some adversarial training works improve the model’s robustness by introducing adversarial perturbations in the embedding space (Zhu et al., 2020), or by incorporating inductive bias to prevent the model from learning spurious correlations (Wang et al., 2021; Madry et al., 2018), as well as methods that increase diversity of ensembled models, leading to globally better robustness (Pang et al., 2019). These methods effectively improve the model’s robustness without significant compromise to its clean accuracy.
In contrast to the empirical methods, certified defenses can provably guarantee model robustness even under sophisticated attackers. One popular certified-defense approach is randomized smoothing (Ye et al., 2020; Zeng et al., 2023), which constructs a set of stochastic ensembles from the input and leverages their statistical properties to provably guarantee robustness. Differential Privacy (DP) defenses such as Lecuyer et al. (2019) are similar to SAFER and RanMASK, achieving certified robustness through the use of DP framework. Note that RanMASK and DP-defenses incur high training overhead. In contrast, Interval Bound Propagation (IBP) methods (Shi et al., 2020b; Jia et al., 2019; Huang et al., 2019) utilizes axis-aligned bounds to confine adversarial examples. Despite their effectiveness and ability to certify robustness, both approaches rely on access to the synonym sets of potential attacks. Randomized-smoothing methods also modify the training phase to reduce the variance of the outputs of the ensemble. Due to the ensemble, they incur significant computational overhead in the inference phase. In contrast, our defense is lightweight and attack-agnostic, and can better control the variance of the outputs by randomizing the latent representations, instead of via randomization in the input space. A summary of the characteristics of AdvFooler and the existing defenses is shown in Table 1.
@@ 3.1 Randomized latent-space defense against adversarial word substitution
Algorithm 1 AdvFooler
Input: a model $f$, an input sentence $x$, scale $\nu$.
@@
6:  return: $z_{L}$
Since the adversary relies on querying the victim model multiple times to find adversarial examples, receiving inconsistent feedback from the model can make it significantly harder for the adversary to find the optimal adversarial perturbations. To fool such an adversary, we introduce stochasticity to the model by randomizing the latent representation of an input $x$. Formally, for $h_{l}$ as the $l-$th layer of the model, we sample an independent noise vector $\epsilon$, which is added to $h_{l}(x)$ as input to the next layer of the model. Without loss of generality, $\epsilon$ is sampled from a Gaussian distribution $\mathcal{N}(0,\Sigma)$ with $\Sigma$ as a diagonal covariance matrix, or $\mathcal{N}(0,\nu I),\nu\in\mathcal{R}$. The detailed algorithm is presented in Algorithm 1. Note that, instead of randomizing the latent representation, the defender can directly randomize the textual input; however, this approach makes it significantly harder to control the corresponding change in the prediction, or to maintain an acceptable level of the model’s performance when applying the defense.
Let $f_{\operatorname{AdvFooler}}$ be the proposed randomized model corresponding to the original $f$. When the variance of injected noise is low, we can assume that small noise slightly changes the output logits but does not shift the prediction of the model.
@@ 3.2 Effect of randomizing latent representations on adversarial attacks
In textual adversarial attacks, the adversary first determines important parts in a sentence $x=\{w_{0},\dots,w_{l}\}$ to reduce the search space by using the importance score, which is defined as the difference in the confidence:
$I_{\tilde{x}}=f_{y}(x)-f_{y}(\tilde{x}),$ (2)
where $f_{y}$ is the logit returned by the model $f$ w.r.t ground-truth label $y$ and $\tilde{x}$ is the perturbed sentence. For example, Li et al. (2018) and Jin et al. (2020) remove a word $w_{i}$ in the sentence $x$, while Li et al. (2020) replace $w_{i}$ by the token $MASK$ to obtain $\tilde{x}=\{w_{0},\dots,w_{i-1},MASK,w_{i+1},\dots,w_{l}\}$. On the other hand, phrase-level attacks transform $x$ by paraphrasing constituents extracted from the syntactic tree Chen et al. (2021) or masking them out Lei et al. (2022). The adversary then selects perturbations with the highest scores. Intuitively, this process emulates estimating the gradient of the importance score in the discrete space.
When introducing randomness to the latent space, the model output is changed. Such changes can mislead the attacker to select ineffective perturbations. For example, the attacker may instead select “less important” words (i.e., those that have lower original importance scores before randomization) to perturb, thus reducing the attack’s success probability Mrkšić et al. (2016); Li et al. (2018).
Theorem 3.1.
If a random vector $v\sim\mathcal{N}(0,\nu I)$ where $\nu$ is small is added to the hidden layer $h$ of the model $f$ which can be decomposed into $f=g\circ h$, the new important score $I^{\text{new}}_{\tilde{x}}$ is a random variable that follows Gaussian distribution $\mathcal{N}(I_{\tilde{x}},\nu(\|\nabla_{h(x)}g_{y}(h(x))\|_{2}+\|\nabla_{h(\tilde{x})}g_{y}(h(\tilde{x}))\|_{2}))$.
Theorem 3.1 (proof is in the Supplementary) states that randomly perturbing the hidden presentation leads to a randomized important score. When $\nu$ is high, the randomized importance scores have high variance, which can cause the attack to select the wrong important words or phrases.
@@ 3.3 Effect of randomizing latent representations on clean accuracy
If randomization induces substantial fluctuation in the model’s output, there is a possibility that the model might predict a wrong label, resulting in a decrease in clean accuracy. Randomized defenses such as SAFER (Ye et al., 2020) and RanMASK (Zeng et al., 2023) introduce randomness to the model in the input space (i.e., randomizing the input tokens). While they are also effective against textual adversarial attacks, the input space of NLP models is discrete. Perturbing discrete inputs can lead to significant changes in the model’s outputs (a fact that we will empirically prove in the next section), which results in a different model’s prediction. To mitigate this problem, existing randomized input defenses rely on randomized smoothing: train the model with perturbed input and apply ensemble at inference to reduce the variance in the model’s output. Nevertheless, these approaches are computationally expensive and require access to the training phase. Conversely, by randomizing the latent representations, AdvFooler induces a significantly smaller variance in the model’s output. Furthermore, AdvFooler gives the defender flexibility to select the suitable induced variance without requiring access to the training process. Specifically, given a pretrained classification model and a small clean test set, the defender can select the noise scale $\nu$ at which the variance causes the clean accuracy drops by a chosen percentage (e.g., in our experiments, it is 1%).